OpenStudy (felisha):

How can I determine whether a system has one solution, no solution, or an infinite number? Also how can I determine a system is a pair of parallel lines, intersecting lines, or one line?

Mehek (mehek14):

Are you asking if you're given 2 equations or a graph?

Mehek (mehek14):

Also, Parallel lines = no solution Intersecting lines = one solution One line = infinite solutions

OpenStudy (felisha):

graphing and 2 equations

Mehek (mehek14):

Graph 

Mehek (mehek14):

If you get 2 equations, you can solve by substitution or elimination

OpenStudy (felisha):

The 2 equations doesn't have an ordered pair so would it be elimination

Mehek (mehek14):

No I'm saying, substitution and/or elimination are the 2 ways to solve if you have : y = 3 + x and y = 5 - 2x you can `substitute` the first equation instead of y in the second equation 3 + x = 5 - 2x